I need help with this problem:
Give the center and radius of the circle x^2+y^2+4x-2y=4 and then scatch its graph.
Thank you very much.

Answer by Edwin McCravy(20086)

x² + y² + 4x - 2y = 4

Rearrange the terms and skip some spaces:

x² + 4x     + y² - 2y     = 4 

Take half the coefficient of x which is 4, get 2
Square 2, get 4.  Add + 4 in the first space on 
the left and also add + 4 to the right side:

x² + 4x + 4 + y² - 2y     = 4 + 4

Take half the coefficient of y which is -2, get -1
Square -1, get +1.  Add + 1 in the second space on
the left and also add + 1 to the right side:

x² + 4x + 4 + y² - 2y + 1  = 4 + 4 + 1

Combine the terms on the right side

x² + 4x + 4 + y² - 2y + 1  = 9

Factor the first three terms on the left as
(x + 2)(x + 2) or (x + 2)²

   (x + 2)² + y² - 2y + 1  = 9

Factor the last three terms on the left as
(y - 1)(y - 1) or (y - 1)²

       (x + 2)² + (y - 1)² = 9

Compare that to

      (x - h)² + (y - k)² = r²

-h = +2 so h = -2
-k = -1 so k = 1
r² = 9  so r = 3

So this is a circle with center (h,k) = (-2,1) 
and radius 3

So we get a compass and stick the sharp point 
on the green point below:



Then open out the compass to a radius of 3 and 
swing it around to draw this circle:



To check it we can see that the graph goes 
through these four points:

(-5,1), (-2,4), (1,1) and (-2,-2)

We check those points to see if they all 
satisfy the original equation:

        x² + y² + 4x - 2y = 4

Checking (-5,1)

5² + (-1)² + 4(-5) - 2(1) = 4

          25 + 1 - 20 - 2 = 4
              26 - 20 - 2 = 4
                    6 - 2 = 4
                        4 = 4

That one does.

Checking (-2,4)

(-2)² + 4² + 4(-2) - 2(4) = 4

           4 + 16 - 8 - 8 = 4
               20 - 8 - 8 = 4
                   12 - 8 = 4
                        4 = 4

That one does.

Checking (1,1)

    1² + 1² + 4(1) - 2(1) = 4
            1 + 1 + 4 - 2 = 4
                2 + 4 - 2 = 4
                    6 - 2 = 4
                        4 = 4
That one does.

Checking (-2,-2)

(-2)² + (-2)² + 4(-2) - 2(-2) = 4

                4 + 4 - 8 + 4 = 4
                    8 - 8 + 4 = 4
                            4 = 4
That one does too. So now we know we have the
right circle.

Edwin